# Scaling Multi-Agent Epistemic Planning through GNN-Derived Heuristics

### v2

@@ 4. Experiments and Evaluation
To conduct our experiments, we developed deep, a modernized re-implementation of the state-of-the-art epistemic planner EFP (Fabiano et al., 2020; Fabiano et al., 2024). deep serves as the primary platform for our evaluation and features a modular design that supports the integration of diverse heuristics, including GNN–based ones. Full implementation details and usage instructions are available in the code repository accompanying this work.
Here we present a comparative analysis of our primary contribution, deep equipped with HFS∗—a depth-augmented variant of HFS detailed in Section 4.1—alongside GNN-based heuristics (denoted as GNN), and deep using Breadth-First Search (denoted as BFS). All executions employed bisimulation-based state reduction and visited-state checks, as introduced by Fabiano et al. (2020). For completeness, we also provide comparisons with the different heuristics implemented in -EFP (Fabiano et al., 2024). All of these heuristics remain available in deep and can be combined with the GNN-based one in a portfolio approach. The solver is already equipped to perform this integration easily as a runtime option. We omit the results of such integration, as the focus of this paper is on the use of ML in epistemic planning, rather than on the planner itself.
While several other MEP solvers exist (Muise et al., 2015; Fabiano et al., 2020; Bolander et al., 2021), we focus our evaluation on BFS to emphasize the impact of incorporating learned heuristics guidance into MEP solving. All experiments were conducted with a timeout of 600 seconds on a 13th Gen Intel(R) Core(TM) i9-13900H CPU with 64 GB of system RAM and an NVIDIA RTX 4070 GPU with 8 GB of VRAM.
The evaluation encompasses several standard benchmarks in the MEP setting. For brevity the definitions of all the parameters used for training, along with detailed descriptions of the domains, are provided in Appendix B and D, respectively.
@@
Our primary evaluation metric is the number of nodes expanded during the search (), which reflects the informativeness of the heuristics. We focus on this measure because GNN currently does not leverage batch computation, making runtime comparisons less meaningful. For completeness, we also report plan length () and solving times in ms (). Complete evaluations are provided in Appendix G.
For aggregate metrics, we report the Interquartile Mean () and IQR standard deviation (-) (Agarwal et al., 2021), focusing on problems solved by all approaches.
4.1. Ablation Study
In this section, we conduct the ablation study. It is important to note that our primary goal is to provide a proof of concept demonstrating the integration of data-driven heuristics learning within MEP. Therefore, we defer a comprehensive ablation study on the GNN architecture to future work, as this would require a significant engineering and computational effort. Here, we focus instead on the data representation and the search strategy. Any further improvement on the GNN architecture will naturally build upon this starting point.
In particular, we compare the three e-state representations introduced in Section 3.2 and the two search strategies, HFS and its variant HFS∗. The latter, inspired by the classical A∗ algorithm444Since the GNN-based heuristics is statistical in nature, its admissibility cannot be guaranteed; hence we do not refer to it as A∗. (Foead et al., 2021; Hart et al., 1968), augments the heuristics value predicted by the GNN with the search depth to form the total score:
@@ 4.1. Ablation Study
Table 1 reports results for the three e-state representations—mask, hash, and map—corresponding respectively to the bitmask, hashing, and symbolic ID formulations. As shown in Table 1, hash achieves the best overall performance. This can be attributed to the intrinsic richness of the e-state representation, which enables effective information retrieval even when part of the original information is discarded. Although the bitmask representation is only slightly outperformed by the hashed one in specific cases, we consider the trade-off offered by hash—lighter training requirements, faster data generation, and reduced inference time—well justifies choosing this as the best encoding.
@@ 4.3. Discussion
GNN demonstrates informative and robust performance across a variety of planning domains, as shown by the promising aggregate metrics in Table 3. Except for the GR domain, the table shows a consistent reduction in the number of explored nodes compared to uninformed search, highlighting the efficiency gains achieved by GNN. The GR domain poses a particular challenge due to its sparse solution density: small heuristics inaccuracies can lead to poor navigation, and similar instances or goals may require substantially different plans, confusing GNN. This motivates the portfolio-based heuristics selection implemented in deep, where other heuristics can compensate when one underperforms. We plan to explore a tighter integration of GNN-based heuristics into this portfolio approach as future work.
Additional evidence of scalability is presented in Table 4, where GNN outperforms the baseline in terms of the number of explored nodes, demonstrating its ability to generalize across domains, including those unseen during training.
Results for remain comparable across methods, suggesting that GNN maintains near-optimal solution quality.
Solving times are similarly close, though GNN incurs a slight overhead due to the lack of batch computation.
Table 5 further compares GNN against individual heuristics within -EFP. GNN performs on par with the best individual heuristics, demonstrating its strength and viability as an alternative approach.
Overall, GNN achieves consistent reductions in explored nodes and offers a scalable, learning-based step toward effective heuristics for multi-agent epistemic planning, where such heuristics are currently limited.
@@ 5. Limitations and Future Directions
While our approach achieves promising experimental results, several limitations remain. First, we acknowledge that our current implementation does not yet achieve competitive runtime performance compared to existing heuristics methods. This limitation is largely due to engineering considerations. A proper integration of CUDA-based computation and batch processing, combined with a search strategy capable of exploiting these features while minimizing memory exchange, would, in fact, significantly improve inference speed. Although we recognize this shortcoming, we emphasize that this work is foundational. Addressing the engineering challenges required for an optimized implementation would constitute a substantial effort in its own right—worthy of dedicated study—and represents an important avenue for future research. Nonetheless, our primary goal here is to provide a proof of concept highlighting the potential of heuristics learning in MEP. This is also why we focus on the number of expanded nodes as our primary evaluation metric.
@@
Finally, an important next step is to integrate our GNN-based heuristics estimates into more advanced and potentially more effective search frameworks, such as MCTS Sutton et al. (1998). We believe this direction holds substantial promise for improving the scalability and adaptability of MEP solvers.
@@ 6. Related Works
Traditionally, planning heuristics are either hand-crafted or derived from structural features of the search space (Russell and Norvig, 2020, Ch. 11). ML-based heuristics offer a scalable alternative, learning meaningful patterns from data (Jiménez et al., 2012; Chen et al., 2024). This is exemplified by systems like AlphaGo (Silver et al., 2016), where learned guidance enables scalable MCTS. Our work builds on these ideas but targets a more structured setting, where planning states are represented as Kripke structures. This introduces challenges, which we address using GNNs to extract semantic features from epistemic states. GNNs have also proven effective in classical planning, where they model relational graphs (Silver et al., 2021), learn numeric heuristics (Borelli et al., 2025), or guide adaptive search (Du et al., 2025).
Recent efforts have also explored using Large Language Models (LLMs) in planning. While LLMs are ineffective as standalone planners (Kambhampati et al., 2024; Pallagani et al., 2024), they can aid heuristics generation (Kambhampati et al., 2024; Corrêa et al., 2025) or domain formalization (Tantakoun et al., 2025). However, due to the structured nature of MEP, we believe GNNs are a more suitable choice. We leave the integration of LLMs in this context to future work.
@@
Most work on MEP has focused on foundational problems such as the investigation of DEL fragments (Burigana et al., 2023b), the definition of action languages (Muise et al., 2015; Baral et al., 2022; Burigana and Fabiano, 2022), and the development of underlying representations (Fabiano et al., 2020; Burigana et al., 2023a). While these are fundamental contributions, this paper pursues a different goal: enabling efficient exploration via informed search.
To our knowledge, only one line of work addresses this challenge, namely Le et al. (2018); Fabiano et al. (2024), which derive heuristics using the planning graph structure. Our approach differs by employing data-driven methods. As mentioned, a future direction is to investigate the interplay between these two types of heuristics either in parallel or in conjunction.
Other recent efforts integrate ML and RL with epistemic planning but simplify epistemic state representations. Engesser et al. (2025) decompose e-states into bounded feature vectors, bridging epistemic logic and reinforcement learning. Similarly, Nunn et al. (2024) use generative models to reason over individual formulae rather than full Kripke structures, enabling different but interesting capabilities.
@@ Appendix C Goal Encoding into e-State Embedding
Moreover, as shown in Bolander et al. (2015), the computational complexity of generating such structures is exponential in complexity with respect to the size of the formula.
@@ Appendix D Domains
Here we present the complete description of the benchmarks used to evaluate our contributions. These have been collected from the literature (Kominis and Geffner, 2015; Huang et al., 2017; Cooper et al., 2019; Fabiano et al., 2020).
@@ Appendix E Forward Pass Neural Distance Regressor
Appendix F Ablation Study
In what follows, we present the results for all the experiments of the ablation study.
@@ Appendix G Experimental Results
-EFP: a planner that also incorporates heuristics, albeit not derived through learning techniques (Fabiano et al., 2024). Following Fabiano et al. (2024), we will use the following to distinguish between the various heuristics used by -EFP

### v3

@@ 4. Experiments and Evaluation
To conduct our experiments, we developed deep, a modernized re-implementation of the epistemic planner EFP (icaps20; DBLP:conf/prima/FabianoPSP24). deep serves as the primary platform for our evaluation and features a modular design that supports the integration of diverse heuristics, including GNN–based ones. Full implementation details and usage instructions are available at https://github.com/FrancescoFabiano/deep.
Here we present a comparative analysis of our primary contribution, deep equipped with HFS∗—a depth-augmented variant of HFS detailed in Section 4.1—alongside GNN-based heuristics (denoted as GNN), and deep using Breadth-First Search (denoted as BFS). All executions employed bisimulation-based state reduction and visited-state checks, as introduced by icaps20. For completeness, we also provide comparisons with the different heuristics implemented in -EFP (DBLP:conf/prima/FabianoPSP24). All of these heuristics remain available in deep and can be combined with the GNN-based one in a portfolio approach. The solver is already equipped to perform this integration easily as a runtime option. We omit the results of such integration, as the focus of this paper is on the use of ML in epistemic planning, rather than on the planner itself.
While several other MEP solvers exist (muise2015planning; icaps20; KR2021-12), we focus our evaluation on BFS to emphasize the impact of incorporating learned heuristics guidance into MEP solving. All experiments were conducted with a timeout of 600 seconds on a 13th-generation Intel® Core™ i9-13900H CPU with 64 GB of system RAM and an NVIDIA RTX 4070 GPU with 8 GB of VRAM.
The evaluation encompasses several standard benchmarks in the MEP setting. For brevity, the definitions of the parameters used for training and the descriptions of the domains, are provided in Appendix B and D, respectively.
@@
Our primary evaluation metric is the number of nodes expanded during the search (), which reflects the informativeness of the heuristics. We focus on this measure because GNN currently does not leverage batch computation, making runtime comparisons less meaningful. For completeness, we also report plan length () and solving times in ms ().
For aggregate metrics, we report the Interquartile Mean () and IQR standard deviation (-) (DBLP:conf/nips/AgarwalSCCB21), focusing on problems solved by all approaches.
4.1. E-State Representation and Search Strategy
In this section, we analyze how the e-state representation and the search strategy affect the overall performance of our approach.
Specifically, we compare the three e-state representations introduced in Section 3.2 and the two search strategies, HFS and its variant HFS∗. The latter, inspired by the classical A∗ algorithm444Since the GNN-based heuristics is statistical in nature, its admissibility cannot be guaranteed; hence we do not refer to it as A∗. (FOEAD2021507; 4082128), augments the heuristics value predicted by the GNN with the search depth to form the total score:
@@ 4.1. E-State Representation and Search Strategy
Table 1 reports results for the three e-state representations—map, hash, and mask—corresponding respectively to the symbolic ID, hashing, and bitmask formulations. As shown in Table 1, hash achieves the best overall performance. This can be attributed to the intrinsic richness of the e-state representation, which enables effective information retrieval even when part of the original information is discarded. Although the bitmask representation is only slightly outperformed by the hashed one in specific cases, we consider the trade-off offered by hash—lighter training requirements, faster data generation, and reduced inference time—well justifies choosing this as the best encoding.
@@ 4.3. Discussion
GNN demonstrates informative and robust performance across a range of planning domains, as evidenced by the aggregate metrics reported in Table 3. With the exception of the GR domain, GNN consistently reduces the number of explored nodes compared to uninformed search, highlighting the efficiency gains enabled by the learned heuristic.
The GR domain presents a particular challenge due to its sparse solution density: even small heuristic inaccuracies can result in poor search guidance, and instances with superficially similar states or goals may require substantially different plans. This variability can mislead GNN and explains its reduced effectiveness in this domain. Importantly, this observation motivates the portfolio-based heuristic selection strategy adopted in deep, where complementary heuristics can compensate when individual ones underperform. A tighter integration of GNN-based heuristics into this portfolio framework is a natural direction for future work.
[…]
Results for are comparable across methods, indicating that GNN preserves near-optimal solution quality. Solving times are likewise similar, although GNN incurs a modest overhead due to the absence of batch computation. A detailed analysis of how batch inference would affect GNN—bringing its per-node overhead closer to that of BFS—is presented in Appendix H.
Table 5 further compares GNN against individual heuristics within -EFP. GNN performs on par with the best individual heuristics, demonstrating its strength and viability as an alternative approach. At the same time, we view GNN-based heuristics as a complementary tool to enhance solver coverage, for example through integration into portfolio-based approaches, rather than as a replacement for existing heuristics.
Overall, GNN achieves consistent reductions in explored nodes and represents a scalable, learning-based step toward effective heuristics for multi-agent epistemic planning, a setting in which heuristic guidance is currently limited.
@@ 5. Limitations and Future Directions
While our approach achieves promising experimental results, several limitations remain. First, we acknowledge that our current implementation does not yet achieve competitive runtime performance compared to existing heuristics methods. This limitation is largely due to engineering considerations. A proper integration of CUDA-based computation and batch processing, combined with a search strategy capable of exploiting these features while minimizing memory exchange, would, in fact, significantly improve inference speed as detailed in Appendix H. Although we recognize this shortcoming, we emphasize that this work is foundational. Addressing the engineering challenges required for an optimized implementation would constitute a substantial effort in its own right—worthy of dedicated study—and represents an important avenue for future research. Nonetheless, our primary goal here is to provide a proof of concept highlighting the potential of heuristics learning in MEP. This is also why we focus on the number of expanded nodes as our primary evaluation metric.
@@
Finally, an important next step is to integrate our GNN-based heuristics estimates into more advanced and potentially more effective search frameworks, such as MCTS sutton1998reinforcement. We believe this direction holds substantial promise for improving the scalability and adaptability of MEP solvers.
@@ 6. Related Works
Traditionally, planning heuristics are either hand-crafted or derived from structural features of the search space (modernApproach, Ch. 11). ML-based heuristics offer a scalable alternative, learning meaningful patterns from data (jimenez2012review; chen2024return). This is exemplified by systems like AlphaGo (silver2016mastering), where learned guidance enables scalable MCTS. Our work builds on these ideas but targets a more structured setting, where planning states are represented as Kripke structures. This introduces challenges, which we address using GNNs to extract semantic features from epistemic states. GNNs have also proven effective in classical planning, where they model relational graphs (silver2021planning), learn numeric heuristics (borelli2025learning), or guide adaptive search (du2025fast).
Recent efforts have also explored using Large Language Models (LLMs) in planning. While LLMs are ineffective as standalone planners (kambhampati2024llmscantplanhelp; DBLP:conf/icaps/PallaganiM0FLMS24), they can aid heuristics generation (kambhampati2024llmscantplanhelp; correa2025classical) or domain formalization (tantakoun-etal-2025-llms). However, due to the structured nature of MEP, we believe GNNs are a more suitable choice. We leave the integration of LLMs in this context to future work.
@@
Most work on MEP has focused on foundational problems such as the investigation of DEL fragments (DBLP:conf/atal/BuriganaFMT23), the definition of action languages (muise2015planning; baral2021action; DBLP:conf/aiia/BuriganaF22a), and the development of underlying representations (icaps20; DBLP:conf/jelia/BuriganaFM23). While these are fundamental contributions, this paper pursues a different goal: enabling efficient exploration via informed search.
To the best of our knowledge, only one line of work addresses this challenge, namely le2018efp; DBLP:conf/prima/FabianoPSP24, which derive heuristics using the planning graph structure. Our approach differs by employing data-driven methods. As mentioned, a future direction is to investigate the interplay between these two types of heuristics either in parallel or in conjunction.
Other recent efforts integrate ML and RL with epistemic planning but simplify epistemic state representations. engesser2025simple decompose e-states into bounded feature vectors, bridging epistemic logic and reinforcement learning. Similarly, nunn2024logic use generative models to reason over individual formulae rather than full Kripke structures, enabling different but interesting capabilities.
@@ Appendix C Goal Encoding into e-State Embedding
Moreover, as shown in bolander2015complexity, the computational complexity of generating such structures is exponential in complexity with respect to the size of the formula.
@@ Appendix D Domains
Here we present the complete description of the benchmarks used to evaluate our contributions. These have been collected from the literature (kominis2015beliefs; huang2017general; cooper:hal-02147986; icaps20).
@@ Appendix E Forward Pass Neural Distance Regressor
Appendix F E-State Representation and Search Strategy
In what follows, we present the results for all the experiments of the e-state representation and search strategy study.
@@ Appendix G Experimental Results
-EFP: a planner that also incorporates heuristics, albeit not derived through learning techniques (DBLP:conf/prima/FabianoPSP24). Following DBLP:conf/prima/FabianoPSP24, we will use the following to distinguish between the various heuristics used by -EFP
@@ G.2. Experimental Setup #2 – Knowledge Transfer
Appendix H Overhead of Single-Query GNN-Based Estimation
Although our current implementation evaluates the GNN heuristic one state at a time, we can estimate the potential impact of batching using aggregate runtime statistics.
From Table 4, BFS requires ms (time per expanded node), whereas the GNN-based configuration requires ms per node. The difference, ms, represents the effective overhead of a single-query GNN evaluation.
Assuming that batching affects only this scoring component, we model the per-node cost under batch size as
where ms denotes the planner-dominated per-node cost. For , this yields ms, matching the observed single-query runtime.
The total projected runtime under batching is then
where is the number of nodes expanded by GNN.
Figure 2 reports the resulting estimates for . The results show that batching rapidly amortizes the GNN overhead: with , the projected runtime decreases from ms to approximately ms, approaching the planner-dominated regime. This supports our choice of expanded nodes as the primary evaluation metric, as heuristic informativeness remains invariant under such engineering optimizations.
Finally, we quantify the relative speed-up induced by batching. Given the projected runtime , we define the speed-up with respect to single-query inference () as
Note that , with diminishing returns as approaches the planner-dominated baseline .